# Can Frontier LLMs Replace Annotators in Biomedical Text Mining? Analyzing Challenges and Exploring Solutions

### v1

@@ 3 Methodology
To address the challenges identified in Section 2.2, we applied a targeted prompt engineering strategy for biomedical text mining tasks. This strategy includes general techniques like dynamic few-shot prompting and some targeted methods like two-step inference for maintaining both reasoning processes alongside structured output, automatic prompt optimization to account for distributional differences between LLM outputs and ground truth labels, and a workflow for dynamically extracting annotation guidelines to provide detailed instructions to LLMs.
The experiments mainly utilized the GPT-4o model (2024-08-06), which was one of the most advanced and widely used models. Additionally, in the later stages of this study, several LLMs supporting test-time compute were released, which are not constrained by missing reasoning steps due to formatting limitations. Based on that, we also evaluated OpenAI o1-mini which is a representative model in this category (4.2).
To ensure that the observed performance improvements were not solely due to advancements in the underlying LLM itself, we also conducted experiments on the LitCovid and BC5CDR-Chemical datasets using historical model consistent with previous studies (see Appendix 7.2). Moreover, we applied model distillation techniques to the HoC and BC5CDR-RE tasks. This involved annotating papers published in April 2024 using LLMs to create augmented datasets, which were subsequently used for fine-tuning BERT-based models.
In the following, we describe these components in detail.
@@ 3.1.1 Dynamic Few-shot Prompting
In-Context Learning (ICL) is a foundational capability of large language models (LLMs), enabling them to solve new tasks by providing instructions or a few examples of the task in the input[22]. Research has shown that the choice of examples used in few-shot prompting significantly affects model performance[50]. Based on that, We implemented a semantic-based dynamic few-shot prompting approach, which is a widely adopted technique across diverse tasks.
As showned in Figure 3 , we embedded each document (a combination of its title and abstract) in the dataset to a high-dimensional embedding and stored these embeddings in a vector database. In the inference phase, the $K$ nearest documents in the embedding space, based on cosine similarity, were retrieved as examples along with their expected outputs. We used OpenAI’s text-embedding-3-small model as the embedding model, Chroma for the vector database, and set $K=5$ for our experiments.
@@ 3.1.2 Strategies for Structured Output
In our study on biomedical multi-label text classification, we adopt JSON as the standard output format. JSON is widely recognized as a structured data format for output generation in large language models (LLMs) and is explicitly supported by APIs such as OpenAI’s. In our implementation for text classification tasks, each key in the JSON structure corresponds to a label, and its associated value indicates whether the label is selected.
However, JSON is not well-suited for Named Entity Recognition (NER) and Relation Extraction (RE) tasks. For NER, the conventional approach involves the BIO format[62], which uses the labels B (Beginning), I (Inside), and O (Outside) to denote whether a token—typically a word or subword—marks the start of an entity, is inside an entity span, or lies outside any entity span. This approach aligns with the pre-LLM paradigm for NER, where the task is framed as sequence labeling: predicting the properties of each individual token. Prior studies such as those by I. Jahan et al., which evaluated LLMs in biomedical NER tasks using the BIO format, showed a very limited performance[27]. LLMs, by design, excel at generating coherent text rather than making predictions for individual tokens. To address this, S. Wang et al. developed a method that uses delimiters to enclose entities, achieving promising results[63]. Building on this insight and considering that contemporary LLM training datasets often include a significant amount of HTML-formatted text, we adopted an HTML tagging scheme for NER tasks. In our approach, entities are marked using tags such as: <Type = "$\{entity\ type\}$"> $\{entity\ mention\}$ </Type>
@@ 3.1.3 Two-step Inference to Maintain Both Reasoning and Structured Outputs
Chain-of-Thought (CoT) is a prompting technique designed to enhance reasoning by incorporating intermediate steps before arriving at the final answer [31]. By breaking down complex problems into smaller, more manageable steps, CoT has been shown to improve the accuracy of foundational model outputs. Frontier large language models (LLMs) are often trained on datasets that include reasoning processes, enabling them to autonomously decompose problems and engage in step-by-step reasoning [64]. As a result, this capability may naturally emerge even in the absence of explicit instructions for problem decomposition in the prompt. However, in biomedical text mining tasks, models are typically required to extract structured results directly from query text in bulk, rather than generating unstructured, difficult-to-parse textual responses.
Some recently released models that support test-time computation do not inherently suffer from this issue. However, for widely used models from the pre-test-time computation era, which offer faster inference (such as GPT-4o), structured constraints often force the model to bypass intermediate reasoning processes. Such omissions can degrade model performance, as demonstrated in Table 2 under “2. Errors due to Inadequate Semantic Comprehension.” Quantitative analyses by Z. Tam et al. further corroborate these findings, highlighting the impact of insufficient reasoning on task outcomes [29].
To address this challenge, we introduce a two-step approach that allows LLMs to first generate complete reasoning and thought processes, and subsequently extract structured answers from them (Figure 3.1.3). For multi-label biomedical text classification tasks requiring JSON-formatted outputs, we inserted an "intermediate steps" schema before the final output in OpenAI’s structured output[65]. For other tasks such as NER and RE which demand alternative output formats, and for the comparison test in LitCovid dataset using the legacy GPT-4 model where the strcutured ouput was not supported in the API, we implemented this functionality independently.
captionWorkflow of the Two-step Inference to Maintain Both Reasoning Step and Structured Outputs
@@ 3.1.4 Automatic Prompt Optimization
In Section 2.2, our analysis reveals a discrepancy between the distribution of labels predicted by LLM and the ground truth labels, and the inability of capturing such implicit features may contribute to their suboptimal performance. To address this issue, we aim to optimize prompts by narrowing the gap between predicted and correct label distributions.
Numerous automatic prompt engineering methods have been developed[37, 36], and we adopt and modify R. Pryzant’s pseudo-“gradient descent” approach which was inspired by gradient descent but the target of optimizatation is the prompt instead of the model parameters[66]. In this method, we identify distributional gap between LLM predictions and ground truth distributions on small batches of data, along with errors in the LLM’s output, to construct a natural language "gradient" that critiques the current prompt. By editing the prompt in the semantic direction opposite to this gradient, we aim to propagate the gradient into the prompt and identify a prompt that optimizes the desired metric (Macro-F1 score) at the dataset level. During the prompt refinement process, multiple candidate prompts are generated, and beam search is employed as the search algorithm. In each iteration, the top k prompts with the highest scores are retained as candidates for the next round, while lower-scoring prompts are discarded.
However, this method is computationally expensive, primarily due to the need for iterative batch inferences. To achieve reliable estimates of distribution differences, we set the batch size to 100. Additionally, implementing this approach for different tasks requires significant coding effort. As a result, our experiments were limited to the LitCovid and BC5CDR-Chemical dataset. For each iteration, we generated three pseudo-gradients along with corresponding candidate prompts during the "propagation" step. In the beam search, the top three candidates were retained in each iteration, and a total of 10 iterations were performed.
@@ 3.2 Instruction Retrieval from Annotation Guideline
Our analysis in Section 2.2 reveals that limitations of LLMs in biomedical text mining arise primarily from their inability to fully comprehend the dataset-specific detailed annotation requirements. At the early stage in our research, we attempted to integrate complete annotation guidelines into the prompts for the NCBI-Disease (NER) and BioRED (RE) datasets to address this issue. However, this approach yielded negligible performance improvements. A plausible explanation is the excessive length of these guidelines, which makes it difficult for LLM to extract the necessary information within such a long context.
Based on that, we explored dynamically extracting relevant entries from the annotation guidelines for each query case. A common approach for such tasks is Retrieval-Augmented Generation (RAG), which integrates retrieval mechanisms with external knowledge bases to provide LLMs with accurate contextual information[67, 68]. While RAG techniques are already implemented in advanced commercial generative models, we could not find a readily available framework suitable for our specific needs. Therefore, we developed a custom rule-based pipeline tailored to these requirements.
For example, in the BioRED (RE) dataset, the annotation guidelines span 16 pages, with around 10 pages dedicated to relationship extraction. Within these, the “relation pairs” section accounts for about 4 pages, and the “relation types” section covers 6 pages. These requirements are further divided into seven subsections based on entity types (e.g., “Disease-chemical,” “Disease-variant & Disease-gene”). Specific rules for edge cases are either grouped in a separate “Others” subsection or scattered across the other subsections. To streamline the process, we segmented and reorganized the guidelines into structured text chunks, as shown in Figure 5 "Knowledge Construction Phase".
During inference, we generate all possible combinations of entities from the entity list and group them based on their types. We process all pairs within the same type in a single query instead of querying each entity pair individually to reduces API costs. Each test query is carried out in three steps (as illustrated in Figure 5 "Inference Phase").
@@
    Relation(Entity Pair) Prediction: We first predict whether a relationship exists between two entities. All candidate entity pairs within the same type are provided as input, and we incorporate the corresponding Relation Pair Requirements into the prompt to instruct the model.
@@
    Relation Type Classification: We then predict the relationship type for the identified entity pairs. Here, the Relation Type Requirements are appended to the prompt to refine the prediction.
@@
    Rule Validation: In the final step, we perform rule validation to verify if the predicted entity pairs and their relationship types require special handling. This step ensures the outputs align with the annotation guidelines and satisfy the necessary constraints.
For the BC5CDR-Chemical/Disease (NER) and NCBI-Disease Named Entity Recognition (NER) datasets, we adopt a similar approach. First, we provide LLM with requirements for identifying and labeling entity mentions to guide the model annotates the mentions accordingly. Then, we integrate instructions for handling edge cases to the prompt and guide LLM to check if the annotations adhere to the specified requirements. The BC5CDR-Chemical/Disease dataset includes two types of entities—Chemical and Disease—which required separate executions of the annotation steps. For the NCBI-Disease dataset, which contains only Disease entities, the steps were executed once. For the BC5CDR (RE) dataset, we simplified the process because the annotation guideline spans less than half a page. We embedded the guideline directly into the prompt, and let the LLM to produce results in a single query.
@@ 3.3 Model Distillation
To explore the potential of large language models (LLMs) in replacing human annotators in the widely adopted “manual annotation–model training–inference” workflow prevalent in real-world applications, we treat the LLM combined with its output parser as a discriminative model. We then conduct a model distillation process from GPT-4o to a BERT-based model. This process leverages data augmentation and weak supervision, which are fundamental strategies in LLM’s model distillation[69].
Specifically, we utilized an LLM to generate labeled data for unannotated research articles, aiming to train a BERT model for the Hallmarks of Cancer (HoC) dataset. We utilized the same search keyword set as the original HoC paper and retrieved PubMed articles published after April 2024 via the NCBI’s API (E-utilities)[70, 71]. This specific timeframe was chosen to minimize the risk of overlap between the retrieved articles and the LLM’s training data. Furthermore, this approach mimics a real-world scenario where novel annotation needs arise, requiring the labeling of newly published articles. After doing the annotation utilizing GPT-4o with two-step inference, we then filtered the retrieved articles to exclude those with no positive labels during the multi-label classification process, and constructed a LLM-annotated corpus of 3,000 articles. This dataset was then used to fine-tune a BioLinkBERT-Large model.
Similarly, for the BC5CDR-RE dataset, we followed the same procedure as in the original study that for each articles in BC5CDR corpus, we retrive all the related articles published after April 2024 by E-utilities. Using the titles and abstracts of these papers, we first performed named entity recognition (NER) to extract entity lists, followed by relation extraction (RE) to identify chemical-disease interactions (CDI) utilizing GPT-4o with two-step inference + instruction retrieval from the annotation guideline. From this set, we built an LLM-augmented corpus of 2361 articles including 5930 CDI records. This corpus was then used to fine-tune the BioLinkBERT-Large model. Negative samples were created by pairing Chemical and Disease entities that did not exhibit CDI relationships, and the number of negative samples was reduced through random sampling to maintain a 2:1 ratio of negative to positive samples.
@@ 4.1 Evaluation of GPT-4o Performance With Proposed Methods
Tables 3, 4, and 5 summarize the results of our experiments for GPT-4o across multi-label text classification, named entity recognition, and relation extraction tasks with targeted prompt engineering. Information about baseline models can be referred in Table 1. Additionally, five out of six failure cases listed in Table LABEL:tab:failures (PMIDs: 24717468, 15485686, 941901, 15036754, 22836123) were successfully addressed using the “Two-step Inference + Instruction Retrieval from Annotation Guideline” approach. The remaining case (PMID: 9400934) was resolved using the “Dynamic 3-shot + Instruction Retrieval from Annotation Guideline” method.
@@
The Dynamic 3-shot approach outperformed the basic instruction method (which used prompts similar to those in prior studies, supplemented with randomly selected 3-shot examples) across all datasets except HoC. This improvement can be attributed to that more contextually relevant examples provided the LLM with additional background knowledge and annotation style about similar queries. However, while Dynamic few-shot does not involve model training or parameter updating, it still leverages manually annotated datasets. Conversely, in the Two-step Inference (Reasoning Step & Structurization Step) approach, few-shot prompting was not utilized due to the impracticality of pre-writing reasoning process for every instance in the training set. Nevertheless, the Two-step Inference method achieved significant performance improvements across all tested datasets. This highlights again that the commonly used structuralization for LLMs in biomedical text mining tasks can severely constrain the reasoning capabilities that frontier LLMs acquired during their training phase.
@@
We also experimented with automatic prompt optimization, iteratively refining prompts by using natural language to describe the distributional differences between predicted and correct labels (resembling a "pseudo-gradient descent" approach). However, the performance gains were modest compared to the huge computational cost, likely because the discrepancy patterns between predicted and correct labels varied across test instances. As a result, a short natural language "gradient" struggled to generalize across the test set. Analyzing the prompts generated during optimization revealed that many included generic but dataset-irrelevant instructions (e.g., rewriting prompts in Markdown, adding overly detailed background introductions) or directly retailed the pseudo-gradient information, such as "the predicted label distribution on the test set was …, while the actual distribution was… Use this information to …" or "Example discrepancy in the last inference:…" Such instructions were clearly unusable by the LLM during inference. Nevertheless, some iterations produced potentially helpful guidance, such as "Be mindful of over-predicting topics such as …. These should only be selected if …". Given the computational cost and limited performance improvement, we cannot recommend this approach for production scenario in biomedical text mining. However, it could serve as a source of inspiration when manually curating prompts.
@@
Instruction Retrieval from Annotation Guidelines generally improved performance, except when combined with Dynamic 3-shot on the BC5CDR (RE) dataset. Its impact was particularly notable for datasets like BioRED, which involve complex entity-relation annotations, or when paired with the Two-step Inference approach. Unfortunately, we could not apply this method to datasets such as LitCovid and HoC due to the absence of detailed annotation guidelines. Across all datasets, combining Two-step Inference with Instruction Retrieval from Annotation Guidelines consistently enhanced GPT-4o’s performance. However, this combination required extracting information and reasoning separately for each entity class, increasing the API usage overhead.
@@
Overall, implementing systematic enhancements targeted for biomedical text mining enabled state-of-the-art LLMs to outperform baseline models on two datasets with minimal reliance on manually annotated data and achieve competitive performance on others. This challenges the prevailing notion in prior studies that LLMs underperform in discriminative tasks within the biomedical domain.

### v2

@@ 3 Methodology
To address the challenges identified in Section 2.2, we applied a series of prompt engineering strategy for biomedical text mining tasks. This strategy includes general techniques like dynamic few-shot prompting, as well as some targeted methods like two-step inference for maintaining both reasoning processes alongside structured output, automatic prompt optimization to account for distributional differences between LLM outputs and ground truth labels, and a workflow for dynamically extracting annotation guidelines to provide detailed instructions to LLMs.
The experiments mainly utilized the GPT-4o model (2024-08-06), which was one of the most advanced and widely used models. Additionally, in the later stages of this study, several LLMs supporting test-time compute were released, which are not constrained by missing reasoning steps due to formatting limitations. Therefore, we also evaluated OpenAI o1-mini which is a representative model in this category (4.2).
To ensure that the observed performance improvements were not solely due to advancements in the underlying LLM itself, we also conducted experiments on the LitCovid and BC5CDR-Chemical datasets using historical model consistent with previous studies (see Appendix 1.2). Moreover, we employed model distillation on the HoC and BC5CDR-RE datasets by annotating recent articles with LLMs to construct augmented datasets, which were then used to fine-tune BERT-based models.
Below, we describe these components in detail.
@@ 3.1.1 Dynamic Few-shot Prompting
In-Context Learning In-Context Learning (ICL) is a foundational capability of LLMs, enabling them to perform new tasks given instructions or a few examples in the input. The choice of examples significantly impacts model performance [50]. We adopted a semantic-based dynamic few-shot prompting approach, which is a widely adopted technique across diverse tasks.
As showned in Figure 3, we embedded each document in the dataset and stored these them in a vector database. In the inference phase, the $K$ nearest documents in the embedding space, based on cosine similarity, were retrieved as examples along with their expected outputs. We used OpenAI’s text-embedding-3-small model for embeddings and Chroma as the vector database.
@@ 3.1.2 Strategies for Structured Output
For biomedical multi-label text classification, we adopt JSON as the standard output format. JSON is widely recognized as a structured data format for output generation in LLM and is explicitly supported by APIs such as OpenAI’s. In our implementation, each key in the JSON structure corresponds to a label, and its associated value indicates whether the label is selected.
However, JSON is not well-suited for NER and RE tasks. For NER, the conventional approach involves the BIO format[62], which uses the labels B (Beginning), I (Inside), and O (Outside) to denote whether a token—typically a word or subword—marks the start of an entity, is inside an entity span, or lies outside any entity span. This approach aligns with the pre-LLM paradigm for NER, where the task is framed as sequence labeling. Prior LLM evaluation using BIO format for NER showed a very limited performance[27]. LLMs, by design, excel at generating coherent text rather than making predictions for individual tokens. To address this, S. Wang et al. developed a method that uses delimiters to enclose entities[63]. Building on this insight and considering that contemporary LLM training datasets often include a significant amount of HTML-formatted text, we adopted an HTML tagging scheme for NER. In our approach, entities are marked using tags such as: <Type = "$\{entity\ type\}$"> $\{entity\ mention\}$ </Type>
@@ 3.1.3 Two-step Inference to Maintain Both Reasoning and Structured Outputs
Chain-of-Thought (CoT) prompting enhances reasoning by explicitly incorporating intermediate steps before providing final answers [31]. Frontier LLM are often trained on datasets that include reasoning processes, enabling them to autonomously decompose problems and engage in step-by-step reasoning even without explicit prompting[64]. However, in biomedical text mining tasks, models are typically required to extract structured results directly from query text in bulk, rather than generating unstructured, difficult-to-parse output. Such omissions can degrade model performance (e.g. Appendix Table S1.b). Quantitative analyses by Z. Tam et al. further corroborate similar findings[29].
Some recently released models that support test-time computation do not inherently suffer from this issue. However, for widely used models released before the test-time computation era, which offer faster inference (such as GPT-4o), structured constraints often force the model to bypass intermediate reasoning processes. To resolve this, we introduce a two-step approach that allows LLMs to first generate complete reasoning processes, and subsequently extract structured answers from them (Figure 4). For multi-label biomedical text classification tasks requiring JSON outputs, we inserted an "intermediate steps" schema before the final output in OpenAI’s structured output API[65]. For NER and RE tasks which demand alternative output formats, and for the experiment in Appendix 1.2, we implemented this functionality independently.
Figure 4: Workflow of the Two-step Inference to Maintain Both Reasoning Step and Structured Outputs
@@ 3.1.4 Automatic Prompt Optimization
In Section 2.2, our analysis reveals a discrepancy between the distribution of labels predicted by LLM and the ground truth labels, and the inability of capturing such implicit features may contribute to their suboptimal performance. To address this, we aim to optimize prompts by narrowing the gap between predicted and correct label distributions.
We adopt and modify R. Pryzant’s pseudo-“gradient descent” approach which was inspired by gradient descent but the target of optimizatation is the prompt instead of the model parameters[66]. This method identifies distributional gaps on small batches of data and constructs a natural-language “gradient” critiquing the current prompt. By editing the prompt semantically opposite to this gradient, we iteratively refine prompts that optimize the dataset-level metric (Macro-F1). Multiple candidate prompts are generated in each iteration, using beam search to select top candidates for subsequent refinement rounds. Given its high computational cost, our experiments were limited to the LitCovid and BC5CDR-Chemical datasets. Detailed descriptions are provided in Appendix 1.3. due to space constraints.
@@ 3.2 Instruction Retrieval from Annotation Guideline
Our analysis in Section 2.2 reveals that the limitations of LLMs in biomedical text mining arise primarily from their insufficient comprehension of detailed, dataset-specific annotation requirements. Initially, we attempted to integrate complete annotation guidelines into the prompts, but yielded negligible improvements, possibly due to the excessive length of the guidelines hindering the LLM’s effective utilization of information within long contexts.
Thus, we explored dynamically extracting relevant entries from the annotation guidelines for each query case. A common approach for such tasks is Retrieval-Augmented Generation (RAG), which integrates retrieval mechanisms with external knowledge bases to provide LLMs with accurate contextual information[67, 68]. While RAG techniques are already implemented in advanced commercial generative models, we could not find a readily available framework suitable for our specific needs. Therefore, we developed a custom rule-based pipeline tailored to these requirements.
For example, in the BioRED (RE) dataset, the annotation guidelines span 16 pages, with around 10 pages dedicated to relationship extraction. Within these, the “relation pairs” section accounts for about 4 pages, and the “relation types” section covers 6 pages. These requirements are further divided into seven subsections based on entity types (e.g., “Disease-chemical,” “Disease-variant”). Specific rules for edge cases are grouped in a separate “Others” subsection. To facilitate retrieval, we reorganized these guidelines into text chunks (Figure 6, Knowledge Construction Phase).
During inference, processing each test query involves three steps (Figure 6, Inference Phase). Alos, instead of querying each candidate entity pairs individually, all candidate pairs from two specific entity types are queried simultaneously to reduce inference costs.
@@
    Relation(Entity Pair) Prediction: Predicts if relationships exist between entity pairs using the relevant Relation Pair Requirements.
@@
    Relation Type Classification: Classifies the relation types of identified pairs using the relevant Relation Type Requirements.
@@
    Rule Validation: Checks whether predicted pairs and types require special handling and conform to guideline constraints, ensuring accuracy.
For other datasets accompanied by annotation guidelines, we adopted a similar workflow. For the BC5CDR-Chemical/Disease (NER) and NCBI-Disease (NER) datasets, we first provided the LLM with simplified requirements on annotating entity mentions, manually summarized from the annotation guidelines, instructing the model to generate preliminary annotations. Subsequently, we incorporated the complete annotation guidelines including examples and edge cases into the prompt, guiding the LLM for detailed rule validation. For the BC5CDR relation extraction (RE) dataset, since its annotation guideline is concise (less than half a page), we directly integrated the guideline into the prompt, enabling the LLM to produce results with a single query.
@@ 3.3 Model Distillation
To explore the potential of LLMs in replacing human annotators in the widely adopted “manual annotation–model training–inference” workflow in real-world applications, we conducted model distillation from GPT-4o to a BERT-based model using data augmentation and weak supervision, which are fundamental strategies in LLM distillation.[69].
For the HoC dataset, we used the original search keywords to retrieve PubMed articles published after April 2024 via NCBI’s E-utilities API[70, 71]. We annotated these articles with GPT-4o using two-step inference, then filtered out articles without positive labels to construct an LLM-annotated corpus of 3,000 articles. This synthetic dataset was then used to fine-tune a BioLinkBERT-Large model.
Similarly, for the BC5CDR-RE dataset, following the same data collection procedure as in the original study, we retrive all related articles published recently. Using the titles and abstracts of these papers, we first performed named entity recognition to extract entity lists, followed by relation extraction to identify chemical-disease interactions (CDI) utilizing GPT-4o with two-step inference + instruction retrieval from the annotation guideline. We created an augmented corpus of 2,361 articles containing 5,930 CDI records, which was then used to fine-tune BioLinkBERT-Large. Negative samples were generated by randomly pairing Chemical-Disease entities without CDI relationships, maintaining a 2:1 negative-to-positive ratio.
The timeframe after April 2024 was selected to minimize overlap between retrieved articles and the GPT-4o’s training data. Furthermore, this approach mimics a real-world scenario where novel annotation needs arise, requiring the labeling of newly published articles.
@@ 4.1 Evaluation of GPT-4o Performance With Proposed Methods
Tables 3, 4, and 5 summarize GPT-4o’s performance across with targeted prompt engineering. Additionally, five out of six failure cases listed in Table LABEL:tab:failures and Appendix Table 1 (PMIDs: 24717468, 15485686, 941901, 15036754, 22836123) were successfully addressed using the “Two-step Inference + Instruction Retrieval from Annotation Guideline” approach. The remaining case (PMID: 9400934) was resolved using the “Dynamic 3-shot + Instruction Retrieval from Annotation Guideline” method.
@@
The Dynamic 3-shot approach outperformed the basic instruction prompts (i.e., prompts similar to those used in prior studies) across all datasets, except for HoC. This can be attributed to that more contextually relevant examples provided the LLM with additional background knowledge and annotation style about similar queries. However, this method still indirectly relies on manually annotated data. Conversely, in the Two-step Inference (Reasoning Step & Structurization Step) approach, few-shot prompting was not utilized due to the impracticality of pre-writing reasoning process for every instance in the training set. Nevertheless, the Two-step Inference method achieved significant performance improvements across all tested datasets. This highlights again that the commonly used structuralization for LLMs in biomedical text mining tasks can severely constrain the reasoning capabilities that frontier LLMs acquired during their training phase.
@@
We also experimented with automatic prompt optimization, iteratively refining prompts by using natural language to describe the distributional differences between predicted and correct labels (resembling a "pseudo-gradient descent" approach). However, the performance gains were modest compared to the huge computational cost, likely because the discrepancy patterns between predicted and correct labels varied across test instances. As a result, a short natural language "gradient" struggled to generalize across the test set. Analyzing the prompts generated during optimization revealed that many included generic but dataset-irrelevant instructions (e.g., rewriting prompts in Markdown, adding overly detailed background introductions) or directly retelled the pseudo-gradient information, such as "the predicted label distribution on the test set was …, while the actual distribution was… Use this information to …"" Such instructions were clearly unusable by the LLM during inference. Nevertheless, some iterations produced potentially helpful guidance, such as "Be mindful of over-predicting topics such as …. These should only be selected if …". Given the computational cost and limited performance improvement, we cannot recommend this approach for production scenario in biomedical text mining. However, it could serve as a source of inspiration when manually curating prompts.
@@
Instruction Retrieval from Annotation Guidelines generally improved performance, except when combined with Dynamic 3-shot on the BC5CDR (RE) dataset. Its impact was particularly notable for datasets like BioRED, which involve complex entity-relation annotations, or when paired with the Two-step Inference approach. Unfortunately, we could not apply this method to datasets such as LitCovid and HoC due to the absence of detailed annotation guidelines.
@@
Overall, implementing systematic enhancements targeted for biomedical text mining enabled GPT-4o to outperform baseline models on two datasets with minimal manual annotation dependency and achieve competitive performance on others. This challenges the prevailing notion in prior studies that LLMs underperform in discriminative tasks within the biomedical domain.